# Optimal Multi-Fidelity Best-Arm Identification

### v1

@@ 1 Introduction
In some applications, like physics, parameter studies, or hyper-parameter optimization, getting a sample from the arm distribution might be expensive since it requires evaluating or training a complex model and is computationally demanding. However, it is often the case that cheaper, less accurate sampling methods are available, for instance, by using a coarser model in the physics study example. The multi-fidelity bandit framework takes such scenarios into account. When choosing an arm, the algorithm also chooses a fidelity, with a trade-off: a higher fidelity gives a more precise observation but has a higher cost. We assume that the algorithm knows both the cost and the maximal bias of the observations from each fidelity. This is also how the knowledge about the fidelity was modeled in prior work [see, e.g., 16, 15, 25, 31]. The goal is then to find the best arm (i.e., the arm with the highest mean at the highest fidelity) with high probability and minimal cost.
@@
After presenting additional related works, in Section 2, we define fixed-confidence best arm identification in multi-fidelity bandits in more mathematical detail and introduce the notations used throughout the paper. Then, Section 3 contains our first contribution: a tight instance-dependent lower bound on the cost complexity of any algorithm expressed with the maximum of a complex function over all possible cost allocations. We also highlight features of that lower bound, like the existence of an optimal fidelity for each arm, which should be chosen exclusively. In Section 4, we propose a computationally efficient procedure for computing gradients of the function featured in the lower bound and describe a gradient-based algorithm whose cost complexity is asymptotically matching the lower bound. Finally, in Section 5, we present the results of numerical experiments which demonstrate the good empirical performance of our new algorithm compared to prior work.
@@ 3.2 Sparsity of the oracle weights: a tight concept of optimal fidelity
We conclude our study of the lower bound by further analyzing the optimal allocation $\bm{\omega}^{*}$. Unlike in the standard best arm identification problem, we did not find an efficient algorithm to compute it, which prevents us from using a Track-and-Stop-like approach [8]. Nevertheless, we will explain in the next section how to efficiently compute the $f_{i,j}$ functions and their gradient. These computations are crucial for our algorithm but also allow us to prove our next result about the possible sparsity of $\bm{\omega}^{*}$. For each arm $a\in[K]$, it is not difficult to show that there must exist some fidelity $m\in[M]$ for which $\omega_{a,m}^{*}>0$ (Lemma B.2). However, as the following result highlights, in most cases, only one fidelity per arm has non-zero weight.
@@
Let $\Delta_{K\times M}^{*}(\bm{\mu})\coloneqq\operatornamewithlimits{argmax}_{\bm{\omega}\in\Delta_{K\times M}}F(\bm{\omega},\bm{\mu})$ and
~$\widetilde{\mathcal{M}}_{\textup{MF}}\coloneqq\left\{\bm{\mu}\in\mathcal{M}_{\mathrm{MF}}^{*}:\exists i\in[K],\exists m_{1},m_{2}\in[M]^{2},\exists\bm{\omega}^{*}\in\Delta^{*}_{K\times M}(\bm{\mu}):\omega^{*}_{i,m_{1}}>0,\omega^{*}_{i,m_{2}}>0\right\}.$
The set ~$\widetilde{\mathcal{M}}_{\textup{MF}}$ is a subset of $\mathbb{R}^{K\times M}$ whose Lebesgue measure is zero.
Theorem 3.2 implies that in almost all multi-fidelity bandits, for any $\bm{\omega}^{*}\in\Delta^{*}_{K\times M}(\bm{\mu})$ and each arm $a\in[K]$, there exists a single fidelity $m^{*}_{a}\in[M]$ for which $\omega^{*}_{a,m^{*}_{a}}>0$ holds. However, we note that this result does not offer an easy way to compute these optimal arm-dependent fidelities. [^3] We provide insights on cases in which it is possible to compute the optimal fidelity in Appendix B.4. Nevertheless, as we shall see in the next section, our algorithm does not actually require identifying these optimal fidelity levels to enjoy optimality guarantees.
Finally, we remark that existing MF-BAI works [25, 31] already proposed notions of optimal, arm-dependent fidelity that the agent should employ to identify the optimal arm $\star$. Nevertheless, as we verify in Appendix B.5, these concepts do not comply with the concept of optimal fidelity that arises from the tight lower bound of Theorem 3.1. In other words, there exist bandit models $\bm{\mu}$ in which following these alternative concepts of optimal fidelity leads to sub-optimal performance.
@@ 6 Conclusions
For fixed-confidence best arm identification in multi-fidelity bandits, we presented a lower bound on the cost complexity and an algorithm with a matching upper bound in the regime of high confidence. The algorithm uses features of the lower bound optimization problem in order to compute its updates efficiently. Unlike prior work, it does not require any assumption or prior knowledge on the bandit instance. Our work also confirmed the existence in most cases of an “optimal fidelity” to explore each arm in the asymptotic regime, and revealed that the intuitive such notions proposed in prior work were inaccurate. Yet, our algorithm does not need to identify these optimal fidelities in order to be asymptotically optimal.
This raises the following question: could the performance of the algorithm be enhanced by exploiting the sparsity pattern? We conjecture that estimating the optimal fidelities accurately may actually be harder than identifying the best arm. However, leveraging some sufficient conditions for $w_{a,m}^{*}=0$ (such as the ones given in Proposition B.6) to eliminate some fidelities and reduce the support of the forced exploration component of the algorithm seems a promising idea. A limitation of our current analysis is that it only provides asymptotic guarantees in the high confidence regime, although our experiments reveal good performance for moderate values of $\delta$. In future work, we will seek a better understanding of the moderate confidence regime [30]. To this end, we may leverage some proof techniques from other works using online optimization that obtain finite-time bounds [4, 5]. On the lower bound side, while $C^{*}(\bm{\mu})$ essentially scales with $K$ due to the sparsity pattern, an interesting open question is whether there is a worse case $\mathcal{O}(KM)$ scaling in the moderate confidence regime, indicating that all fidelities do need to be explored at least a constant amount of times.
@@ B.2 Comparison with existing lower bound
In this section, we provide a formal proof on the sparsity of the optimal oracle allocation $\bm{\omega}^{*}$. The proofs given in this section rely on results that are explained in Appendix C.1.
At this point, in order to prove Theorem 3.2, we first introduce some intermediate results that will be used in the proving the theorem. Specifically, we begin by showing that, for each arm $a$, there always exists a fidelity $m$ such that $\omega^{*}_{a,m}>0$ holds.
@@
We now continue by providing necessary conditions that characterize some key properties of the oracle weights $\bm{\omega}$, which follows from the expression of the gradient of $f_{\star,a}(\bm{\omega},\bm{\mu})$ with respect to $\bm{\omega}$ for $\bm{\mu}\in\mathcal{M}_{\mathrm{MF}}$ (Lemma C.3).
@@
Consider $\bm{\mu}\in\mathcal{M}$ and $\bm{\omega}^{*}\in\Delta_{K\times M}(\bm{\mu})$. Then, for all $a\neq\star$ the following conditions holds:
@@
which, is a concave optimization problem with a non-empty feasible region. Therefore, we can apply the KKT conditions to study the properties of each local optimal point $\bm{\omega}^{*}$ for which the sub-derivatives exist, i.e., from Theorem 4.3 and Corollary C.2, the ones for which the following condition hold: [^9] We notice that a global optimum point clearly satisfies this condition.
@@
At this point, fix any arm $a$ that attains the minimum in Equation (14). Then, from the KKT conditions, we obtain the following system of inequalities:
$\begin{cases}-\frac{\partial}{\omega_{a,m}}f_{\star,a}(\bm{\omega},\bm{\mu})+c-b_{a,m}=0&\text{ }\forall m\in[M]\\ -\frac{\partial}{\omega_{\star,m}}f_{\star,a}(\bm{\omega},\bm{\mu})=0&\text{ }\forall m\in[M]\\ b_{i,m}\omega^{*}_{i,m}=0&\text{ }\forall i\in\{\star,a\},\forall m\in[M]\\ b_{i,m}\geq 0&\text{ }\forall i\in[K],\forall m\in[M]\\ \omega_{i,m}^{*}\geq 0&\text{ }\forall i\in[K],\forall m\in[M]\\ \sum_{i,m}\omega^{*}_{i,m}=1\\ \end{cases}.$
@@
Finally, to conclude the proof, it is sufficient to iterate these arguments for all $a\neq\star$. Indeed, from Lemma B.3, we know that all sub-optimal arms will attain the minimum in Equation (14) at a global optimum $\bm{\omega}^{*}$. ∎
At this point we are ready to prove our main result.
@@
Let us introduce some additional notation. Consider a subset of arm-fidelity pairs $\mathcal{X}\subseteq[K]\times[M]$, and define $\mathcal{G}(\mathcal{X})\subseteq\mathcal{M}$ as the subset of multi-fidelity bandit models $\bm{\mu}$ for which there exists $\bm{\omega}^{*}\in\Delta^{*}_{K\times M}(\bm{\mu})$ such that, for all $(i,m)\in\mathcal{X}$, $\omega^{*}_{i,m}>0$ holds.
@@
~$\widetilde{\mathcal{M}}_{\textup{MF}}\subseteq\bigcup_{i=1}^{K}\bigcup_{m_{1},m_{2}\in[M]^{2}}\mathcal{G}(\left\{(i,m_{1}),(i,m_{2})\right\})\coloneqq\mathcal{Y}.$

### v2

@@ 1 Introduction
In some applications, like physics, parameter studies, or hyper-parameter optimization, getting a sample from the arm distribution might be expensive since it requires evaluating or training a complex model and is computationally demanding. However, it is often the case that cheaper, less accurate sampling methods are available, for instance, by using a coarser model in the physics study example. The multi-fidelity bandit framework takes such scenarios into account. When choosing an arm, the algorithm also chooses a fidelity, with a trade-off: a higher fidelity gives a more precise observation but has a higher cost. We assume that the algorithm knows both the cost and the maximal bias of the observations from each fidelity. This is also how the knowledge about the fidelity was modeled in prior work [see, e.g., 16, 15, 26, 32]. The goal is then to find the best arm (i.e., the arm with the highest mean at the highest fidelity) with high probability and minimal cost.
@@
After presenting additional related works, in Section 2, we define fixed-confidence best arm identification in multi-fidelity bandits in more mathematical detail and introduce the notations used throughout the paper. Then, Section 3 contains our first contribution: a tight instance-dependent lower bound on the cost complexity of any algorithm expressed with the maximum of a complex function over all possible cost allocations. We also study features of that lower bound, like the existence of an optimal fidelity for each arm in $2\times M$ bandits, which should be chosen exclusively. In Section 4, we propose a computationally efficient procedure for computing gradients of the function featured in the lower bound and describe a gradient-based algorithm whose cost complexity is asymptotically matching the lower bound. Finally, in Section 5, we present the results of numerical experiments which demonstrate the good empirical performance of our new algorithm compared to prior work.
@@ 3.2 Sparsity of the oracle weights: a tight concept of optimal fidelity
We conclude our study of the lower bound by further analyzing the optimal allocation $\bm{\omega}^{*}$. Unlike in the standard best arm identification problem, we did not find an efficient algorithm to compute it, which prevents us from using a Track-and-Stop-like approach [8]. Nevertheless, we will explain in the next section how to efficiently compute the $f_{i,j}$ functions and their gradient. These computations are crucial for our algorithm but also allow us to prove our next result about the possible sparsity of $\bm{\omega}^{*}$ in $2\times M$ bandits. For each arm $a\in[K]$, it is not difficult to show that there must exist some fidelity $m\in[M]$ for which $\omega_{a,m}^{*}>0$ (Lemma B.2). However, as the following result highlights, in most cases of $2\times M$ bandits, only one fidelity per arm has non-zero weight.
@@
Let $K=2$, $\Delta_{2\times M}^{*}(\bm{\mu})\coloneqq\operatornamewithlimits{argmax}_{\bm{\omega}\in\Delta_{2\times M}}F(\bm{\omega},\bm{\mu})$ and
~$\widetilde{\mathcal{M}}_{\textup{MF}}\coloneqq\left\{\bm{\mu}\in\mathcal{M}_{\mathrm{MF}}^{*}:\exists i\in[2],\exists m_{1},m_{2}\in[M]^{2},\exists\bm{\omega}^{*}\in\Delta^{*}_{2\times M}(\bm{\mu}):\omega^{*}_{i,m_{1}}>0,\omega^{*}_{i,m_{2}}>0\right\}.$
The set ~$\widetilde{\mathcal{M}}_{\textup{MF}}$ is a subset of $\mathbb{R}^{2\times M}$ whose Lebesgue measure is zero.
Theorem 3.2 implies that in almost all $2\times M$ multi-fidelity bandits, for any $\bm{\omega}^{*}\in\Delta^{*}_{2\times M}(\bm{\mu})$ and each arm $a\in[2]$, there exists a single fidelity $m^{*}_{a}\in[M]$ for which $\omega^{*}_{a,m^{*}_{a}}>0$ holds. However, we note that this result does not offer an easy way to compute these optimal arm-dependent fidelities.  [^3] We provide insights on cases in which it is possible to compute the optimal fidelity in Appendix B.4.
We remark that existing MF-BAI works [26, 32] already proposed notions of optimal, arm-dependent fidelity that the agent should employ to identify the optimal arm $\star$. Nevertheless, as we verify in Appendix B.5, these concepts do not comply with the concept of optimal fidelity that arises from the tight lower bound of Theorem 3.1. In other words, there exist bandit models $\bm{\mu}$ in which following these alternative concepts of optimal fidelity leads to sub-optimal performance.
@@ 6 Conclusions
For fixed-confidence best arm identification in multi-fidelity bandits, we presented a lower bound on the cost complexity and an algorithm with a matching upper bound in the regime of high confidence. The algorithm uses features of the lower bound optimization problem in order to compute its updates efficiently. Unlike prior work, it does not require any assumption or prior knowledge on the bandit instance. Our work also confirmed the existence in most $2\times M$ bandits of an “optimal fidelity” to explore each arm in the asymptotic regime, and revealed that the intuitive such notions proposed in prior work were inaccurate.
This raises the following questions: (i) does the sparsity pattern hold for $K\times M$ bandits, and, (ii) could the performance of the algorithm be eventually enhanced by exploiting this pattern? We conjecture that the estimating the optimal fidelities accurately may actually be harder than identifying the best arm, even in $2\times M$ bandits. However, leveraging some sufficient conditions for $w_{a,m}^{*}=0$ (such as the ones given in Proposition B.6) to eliminate some fidelities and reduce the support of the forced exploration component of the algorithm seems a promising idea. A limitation of our current analysis is that it only provides asymptotic guarantees in the high confidence regime, although our experiments reveal good performance for moderate values of $\delta$. In future work, we will seek a better understanding of the moderate confidence regime [31]. To this end, we may leverage some proof techniques from other works using online optimization that obtain finite-time bounds [4, 5]. On the lower bound side, under the assumption that the sparsity pattern holds for $K\times M$ bandits, $C^{*}(\bm{\mu})$ would essentially scales with $K$ due to this pattern. An interesting question is whether there is a worse case $\mathcal{O}(KM)$ scaling in the moderate confidence regime, indicating that all fidelities do need to be explored at least a constant amount of times.
@@ B.2 Comparison with existing lower bound
In this section, we provide a formal proof on the sparsity of the optimal oracle allocation $\bm{\omega}^{*}$ for $2\times M$ bandits. The proofs given in this section rely on results that are explained in Appendix C.1.
At this point, in order to prove Theorem 3.2, we first introduce some intermediate results that will be used in the proving the theorem. Specifically, we begin by showing that, for each arm $a$, there always exists a fidelity $m$ such that $\omega^{*}_{a,m}>0$ holds. Note that the following Lemma holds for any $K$.
@@
We now continue by providing necessary conditions that characterize some key properties of the oracle weights $\bm{\omega}$, which follows from the expression of the gradient of $f_{\star,a}(\bm{\omega},\bm{\mu})$ with respect to $\bm{\omega}$ for $\bm{\mu}\in\mathcal{M}_{\mathrm{MF}}$ (Lemma C.3). The following result is limited to $2\times M$ bandits.
@@
Consider $\bm{\mu}\in\mathcal{M}$ and $\bm{\omega}^{*}\in\Delta_{2\times M}(\bm{\mu})$. Then, for $a\neq\star$, the following holds:
@@
which, is a concave optimization problem with a non-empty feasible region. Therefore, we can apply the KKT conditions to study the properties of each local optimal point $\bm{\omega}^{*}$ for which the derivatives exist, i.e., from Theorem 4.3 and Corollary C.2, the ones for which the following condition hold: [^9] We notice that a global optimum point clearly satisfies this condition.
@@
We note that this step explicitely uses the fact that $K=2$. In this case, indeed, $F(\bm{\mu},\cdot)$ is differentiable as the min over the different arms is attained at $a\neq\star$. Consequently, $F(\bm{\mu},\bm{\mu})$ rewrites as $f_{\star,a}(\bm{\omega},\bm{\mu})$. Then, from the KKT conditions, we obtain the following system of inequalities:
$\begin{cases}-\frac{\partial}{\omega_{a,m}}f_{\star,a}(\bm{\omega},\bm{\mu})+c-b_{a,m}=0&\text{ }\forall m\in[M]\\ -\frac{\partial}{\omega_{\star,m}}f_{\star,a}(\bm{\omega},\bm{\mu})+c-b_{\star,m}=0&\text{ }\forall m\in[M]\\ b_{i,m}\omega^{*}_{i,m}=0&\text{ }\forall i\in\{\star,a\},\forall m\in[M]\\ b_{i,m}\geq 0&\text{ }\forall i\in[K],\forall m\in[M]\\ \omega_{i,m}^{*}\geq 0&\text{ }\forall i\in[K],\forall m\in[M]\\ \sum_{i,m}\omega^{*}_{i,m}=1\\ \end{cases}.$
@@
$\frac{d^{+}(\mu_{a,m_{1}},\eta^{*}_{\star,a}(\bm{\omega}^{*})-\xi_{m_{1}})}{\lambda_{m_{1}}}=\frac{d^{-}(\mu_{\star,m_{2}},\eta^{*}_{\star,a}(\bm{\omega}^{*})+\xi_{m_{2}})}{\lambda_{m_{2}}}~{}\forall m_{1},m_{2}:\omega^{*}_{a,m_{1}},\omega^{*}_{\star,m_{2}}>0,$
thus concluding the proof. ∎
At this point we are ready to prove Theorem 3.2.
@@
Let us introduce some additional notation. Consider a subset of arm-fidelity pairs $\mathcal{X}\subseteq[2]\times[M]$, and define $\mathcal{G}(\mathcal{X})\subseteq\mathcal{M}$ as the subset of multi-fidelity bandit models $\bm{\mu}$ for which there exists $\bm{\omega}^{*}\in\Delta^{*}_{2\times M}(\bm{\mu})$ such that, for all $(i,m)\in\mathcal{X}$, $\omega^{*}_{i,m}>0$ holds.
@@
~$\widetilde{\mathcal{M}}_{\textup{MF}}\subseteq\bigcup_{i=1}^{2}\bigcup_{m_{1},m_{2}\in[M]^{2}}\mathcal{G}(\left\{(i,m_{1}),(i,m_{2})\right\})\coloneqq\mathcal{Y}.$